OpenStudy (anonymous):

Does actually 0.9999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999999....... is not 1

OpenStudy (anonymous):

if \[\frac{1}{1-0.9999999...}=infinity\] im i right?

OpenStudy (anonymous):

It's easy to show that 1 = 0.999.... to infinity

OpenStudy (anonymous):

it's easy too, to show \[\sqrt(x^{2})=x\] not absval x

OpenStudy (anonymous):

Sarcasm? :) If you let x = 0.999... Then 10x = 9.999... 10x - x = 9.999... - x 10x - x = 9.999... - 0.999... 9x = 9 x = 1 QED :)

OpenStudy (raden_zaikaria):



OpenStudy (anonymous):

what ??

OpenStudy (raden_zaikaria):